OpenStudy (anonymous):

Solve for c. ab - c = 9 A. -ab- 9 B. 9 +ab C. ab- 9

OpenStudy (anonymous):

ab-c=9 so substract 'ab' both sides: ab-c-ab=9-ab -c=9-ab multiply both sides by '-1': c=-9+ab or ab-9 so C